OpenStudy (anonymous):

what are the complex roots of x cubed +8 =0

OpenStudy (anonymous):

factor it using the sum of cubes... the complex roots comes from the quadratic factor....

OpenStudy (anonymous):

real cubed root is -2 other two are equally spaced about the circle of radius 2 centered at the origin

OpenStudy (anonymous):

or what dpalnc said x^3+8=0 (x+2)(x^2-2x+4)=0 use the quadratic formula for the second two

OpenStudy (anonymous):

sure it does....